# Suppression of the superfluid Kelvin-Helmholtz instability due to massive vortex cores, friction and confinement

### v1

@@ 1 Introduction
The Kelvin-Helmholtz instability consists in the exponential amplification of infinitesimal fluctuations occurring at the interface between two fluid layers having a relative velocity $\Delta v$ [1, 2]. This process leads to the breakdown of the laminar flow and to the onset of vortical structures [3, 4, 5]. This hydrodynamic instability, well known in the classical context and often considered a precursor of turbulence [6, 7, 8, 9], is observed in various natural phenomena, including cloud formation [10, 11], mixing of oceanic currents [12, 13], and even astrophysical systems [14], where it is suspected to be a trigger mechanism for pulsar glitches [15]. The growth rate of this instability is $\sigma_{c}(q)=q\Delta v/2$ [16, 5], $q$ being the wavenumber of the perturbation. The kinematic viscosity $\mu$ introduces an ultra-violet cutoff $q^{*}\propto\Delta v/\mu$ that halts the runaway behaviour for small wavelengths (i.e. large $q$). The maximum growth rate then scales as $\sigma_{c}^{*}=\sigma_{c}(q^{*})\propto\Delta v^{2}$.
@@
Yet, many open questions remain about the detailed physical processes governing the breakdown of regular vortex arrays, especially in fermionic superfluids, where the motion of quantized vortices is affected by the interactions with the gas of elementary excitations and by the presence of Andreev bound states localized at the vortex cores[47, 48, 49, 50, 51]. This is the case, for instance, of the LENS experiment [46], where an instability growth rate $\sim 3$ times smaller than the one predicted by current theoretical models has been measured. They consider subsonic velocity differences where the Kelvin-Helmholtz instability is the dominant mechanism and its suppression cannot be ascribed to any coupling with acoustic excitations [44]. Given that, the observed mismatch has not received a theoretical explanation yet, hence it calls for further studies.
@@
The structure of the manuscript is the following: in Sec. 2, we focus on the impact of a non-zero vortex-core inertial mass on the properties of regular vortex rows and show its rather general stabilizing effect. This analysis is motivated by the fact that, in many real superfluid systems, vortex cores are often filled, either accidentally or deliberately, by massive particles that provide the topological excitations themselves with a non-zero inertial mass. Typical examples include tracer atoms in superfluid ${}^{4}\mathrm{He}$ [54, 55], quasiparticle bound states in fermionic superfluids even at zero temperature [48, 49, 50, 51], thermal atoms in atomic BECs [56], and atoms of a different species in two-component BECs [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]. Moreover, we introduce dissipation into our analysis to make our point-vortex model as accurate as possible. It turns out that dissipative processes indeed have a stabilizing effect on vortex arrays.
@@ 2 Vortex rows
When set into rotation, superfluids can host elementary excitations in the form of quantized vortices. The superfluid density is depleted in correspondence of the vortex cores, while the superfluid flow swirls around them. In quasi-two-dimensional (2D) configurations, the additional degrees of freedom associated to vortex-line bending become too high-lying in energy, and hence freeze out. As a result, superfluid vortices, despite their inherently quantum nature, can be effectively modeled as classical point-like particles, simplifying the complex dynamics associated with their motion. This modeling approach proves particularly valuable when considering the interactions among multiple vortices. According to the principle of superposition of potential flows, the instantaneous velocity of each quantum vortex corresponds to the vector sum of the velocities induced by all other vortices within the system. In the present work we employ such a method since we deal with quasi-2D superfluids. A straight array of equally spaced vortices, the so called “vortex row" (schematically represented in Fig. 1) constitutes a stationary, but unstable, configuration. As is well known [43], in fact, any perturbation of this regular arrangement is amplified with a characteristic rate
@@ 2.1 Instability of a massive vortex row
As mentioned in the Introduction, in many real superfluid systems quantum vortices are often filled, either deliberately or accidentally, by massive particles, which provide them with an effective inertial mass. The ensuing kinetic-energy term, often overlooked by well-established theoretical models [52, 53, 43], can be easily introduced within the Lagrangian description of superfluid vortex dynamics [60, 62]. To capture the impact of a finite core inertial mass on the superfluid Kelvin-Helmholtz instability, we start in this section by computing the maximum instability growth rate relevant to a massive vortex row. Later in Sec. 4, then, we will develop a similar computation for a massive vortex necklace inside a planar annulus.
@@
as it constitutes the maximum instability growth rate. The latter quantity is shown in Fig. 2 as a function of the number of vortices $N_{v}$ contained in a length $L$, i.e. $N_{v}=L/a$. In the limit of massless cores we recover the result of Eq. (2), i.e., $\sigma^{*}\propto a^{-2}\propto(N_{v}/L)^{2}=(\Delta v/\kappa)^{2}$, where we recall that $\Delta v$ is the velocity difference across the vortex row. Such a quadratic dependence on $\Delta v$ is also found in the analysis of the classical Kelvin-Helmholtz instability, and keeps holding for small number $N_{v}$ of massive vortices. Remarkably, however, for large $N_{v}$ (i.e., small $a$) massive cores feature a maximum instability growth rate which is linear in $N_{v}$:
@@ 2.3 Dissipation-induced suppression of the instability
where $\bm{v}_{s,j}$ is the superfluid velocity in the neighborhood of the vortex [73, 75], and the longitudinal and transverse friction coefficients $\alpha$ and $\alpha^{\prime}$ are both typically positive. In this framework, the maximum instability growth rate (18) reads
@@ 3.1 Point-vortex model
$\mathcal{L}_{a}=\sum_{j=1}^{N_{v}}\left\{\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\dot{\theta}_{j}-\Phi_{j}-\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right\},$ (21)
@@
    $\mathcal{H}_{a}=\sum_{j=1}^{N_{v}}\left(\Phi_{j}+\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right)$ (24)
@@ 3.3 Instability of a massless vortex necklace
$\mathcal{L}_{a}^{\prime}=\sum_{j=1}^{N_{v}}\left\{\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\left(\dot{\theta}_{j}^{\prime}+\Omega\right)-\Phi_{j}-\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right\}$ (31)
@@
$\mathcal{H}_{a}\qquad\to\qquad\mathcal{H}_{a}^{\prime}=\sum_{j=1}^{N_{v}}\left(\Phi_{j}+\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right)-\Omega\sum_{j=1}^{N_{v}}\pi\hbar n_{a}(R_{2}^{2}-r_{j}^{2}).$ (32)
@@
In all cases, the most unstable mode is the one associated to $q^{*}=q_{j=N_{v}/2}=\pi/d_{v}$, where $d_{v}\approx 2\pi r_{0}/N_{v}$ is the intervortex distance, while the mode $q=0$ is always stable [i.e. $\sigma(q=0)=0$]. The latter property is associated to the conservation of the total angular momentum, which follows from system rotational invariance. For $N_{v}\to+\infty$, the points collapse on Eq. (1), which was derived for straight vortex rows in Ref. [43]. The structure of the normal modes for $N_{v}=6$ is illustrated in Fig. 8.
@@ 4.1 Hamiltonian description
($\mathcal{L}\left(\left\{r_{j},\theta_{j}\right\},\left\{\dot{r}_{j},\dot{\theta}_{j}\right\}\right)=\sum_{j=1}^{N_{v}}\left\{\frac{M_{c}}{2}\left(\dot{r}_{j}^{2}+r_{j}^{2}\dot{\theta}_{j}^{2}\right)+\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\dot{\theta}_{j}-\Phi_{j}-\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right\}.$ (40)
@@
$\mathcal{H}=\sum_{j=1}^{N_{v}}\left\{\frac{p_{r_{j}}^{2}}{2M_{c}}+\frac{\left[p_{\theta_{j}}-\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\right]^{2}}{2M_{c}r_{j}^{2}}+\Phi_{j}+\sum_{k=1}^{N_{v}}{}^{\prime}V_{jk}\right\}$ (41)
@@ 4.2 Core-mass-induced suppression of the instability
This relation, illustrated in Fig. 11 as a function of the number of vortices (solid lines), well captures the results of the full numerical liner-stability analysis (dots) for a vortex necklace in an annular domain.
@@ 5 Comparison with experiments
Unfortunately, it is not possible to rigorously perform the linear-stability analysis for a (either massless or massive) vortex necklace subject to frictional forces. The reason is that, in the presence of dissipation, regular vortex necklaces (see Sec. 3.2 and Sec. 4.1) no longer constitute stationary solutions of the associated dynamical systems, a circumstance which prevents the application of the standard linear-stability-analysis machinery described in Sec. 3.3 and Sec. 4.2. However, the problem of introducing dissipation into model (40) can be bypassed by adapting Eqs. (16)-(18), obtained for a vortex row, to the case of a vortex necklace. This can be done upon the substitution $a=2\pi r_{0}/N_{v}$ and, on the basis of what was discussed in Sec. 4.2 (and illustrated in Fig. 11), the resulting equations are expected to well capture the properties of the actual vortex necklace.
@@ 6 Conclusions
In conclusion, we have elucidated the role played by confinement, vortex mass and mutual friction on the dynamical instability of many-vortex systems. Further studies are needed to understand the impact of other effects which may be relevant in this context, such as non-zero temperatures or the coupling of vortices to sound.
@@ Acknowledgements
A. R. received funding from the European Union’s Horizon research and innovation programme under the Marie Skłodowska-Curie grant agreement Vortexons (no. 101062887). P. M. and A. R. acknowledge support by the Spanish Ministerio de Ciencia e Innovación (MCIN/AEI/10.13039/501100011033, grant PID2020-113565GB-C21), and by the Generalitat de Catalunya (grant 2021 SGR 01411). P. M. further acknowledges support by the ICREA Academia program. M. C, A. R. and P. M. would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and the LabEx CARMIN (ANR-10-LABX-59-01) for their support.

### v2

@@ 1 Introduction
The Kelvin-Helmholtz instability consists in the exponential amplification of infinitesimal fluctuations occurring at the interface between two fluid layers having a relative velocity $\Delta v$ [1, 2]. This process leads to the breakdown of the laminar flow and to the onset of vortical structures [3, 4, 5]. This hydrodynamic instability, well known in the classical context and often considered a precursor of turbulence [6, 7, 8, 9], is observed in various natural phenomena, including cloud formation [10, 11], mixing of oceanic currents [12, 13], and even astrophysical systems [14], where it is suspected to be a trigger mechanism for pulsar glitches [15]. The growth rate of this instability is $\sigma_{c}(q)=q\Delta v/2$ [16, 5], $q$ being the wavenumber of the perturbation. In a realistic description, the two uniform flows are connected by a thin shear-layer that smooths out the velocity profile at the interface. It also introduces an ultra-violet cut-off $q^{*}$ which halts the runaway behaviour for small wavelengths (i.e. large $q$) of the instability growth rate [5]. In real classical fluids with viscosity $\mu>0$, one has $q^{*}\propto\Delta v/\mu$ and the maximum growth rate scales as $\sigma_{c}^{*}=\sigma_{c}(q^{*})\propto\Delta v^{2}$.
@@
Yet, many open questions remain about the detailed physical processes governing the breakdown of regular vortex arrays. In Bose superfluids, the presence of kelvon bound states localized at the vortex core (also at zero temperature) is responsible for a non-zero vortex mass which influences the vortex motion [47, 48, 49, 50]. Similarly, in fermionic superfluids the motion of quantized vortices is affected by the interactions with the gas of elementary excitations and by the presence of Andreev bound states localized at the vortex cores [51, 52, 53, 54, 55]. This is the case, for instance, of the LENS experiment [46], where an instability growth rate $\sim 3$ times smaller than the one predicted by current theoretical models has been measured. They consider subsonic velocity differences where the Kelvin-Helmholtz instability is the dominant mechanism and its suppression cannot be ascribed to any coupling with acoustic excitations [44]. Given that, the observed mismatch has not received a theoretical explanation yet, hence it calls for further studies.
@@
The structure of the manuscript is the following: in Sec. 2, we focus on the impact of a non-zero vortex-core inertial mass on the properties of regular vortex rows and show its rather general stabilizing effect. This analysis is motivated by the fact that, in many real superfluid systems, vortex cores are often filled, either accidentally or deliberately, by massive particles that provide the topological excitations themselves with a non-zero inertial mass. Typical examples include tracer atoms in superfluid ${}^{4}\mathrm{He}$ [58, 59], quasiparticle bound states both in fermionic [52, 53, 50, 54, 55] and bosonic superfluids [50] even at zero temperature, thermal atoms in atomic BECs [60], and atoms of a different species in two-component BECs [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72]. Moreover, we introduce dissipation into our analysis to make our point-vortex model as accurate as possible. It turns out that dissipative processes indeed have a stabilizing effect on vortex arrays.
@@ 2 Vortex rows
When set into rotation, superfluids can host elementary excitations in the form of quantized vortices. The superfluid density is depleted in correspondence of the vortex cores, while the superfluid flow swirls around them. In quasi-two-dimensional (2D) configurations, the additional degrees of freedom associated to vortex-line bending become too high-lying in energy, and hence freeze out. When finite-compressibility effects can be neglected, and given their quantized vorticity, superfluid vortices can be effectively modeled as classical point-like particles (or as classical filaments in three-dimensional systems), simplifying the complex dynamics associated with their motion. This modeling approach proves particularly valuable when considering the interactions among multiple vortices. According to the principle of superposition of potential flows, the instantaneous velocity of each quantum vortex corresponds to the vector sum of the velocities induced by all other vortices within the system. In the present work we employ such a method since we deal with quasi-2D superfluids. A straight array of equally spaced vortices, the so called “vortex row" (schematically represented in Fig. 1) constitutes a stationary, but unstable, configuration. As is well known [43], in fact, any perturbation of this regular arrangement is amplified with a characteristic rate
@@ 2.1 Instability of a massive vortex row
As mentioned in the Introduction, in many real superfluid systems quantum vortices are often filled, either deliberately or accidentally, by massive particles, which provide them with an effective inertial mass. Interestingly, vortex mass can be an intrinsic property of a large class of quantum vortices, as it originates from kelvon modes or quasi-particle bound states localized in the vortex core [50]. The ensuing kinetic-energy term, often overlooked by well-established theoretical models [56, 57, 43], can be easily introduced within the Lagrangian description of superfluid vortex dynamics [64, 66]. To capture the impact of a finite core inertial mass on the superfluid Kelvin-Helmholtz instability, we start in this section by computing the maximum instability growth rate relevant to a massive vortex row. Later in Sec. 4, then, we will develop a similar computation for a massive vortex necklace inside a planar annulus.
@@
as it constitutes the maximum instability growth rate. The latter quantity is shown in Fig. 2 as a function of the number of vortices $N_{v}$ contained in a length $L$, i.e. $N_{v}=L/a$. In the limit of massless cores ($M_{c}\to 0$) we recover the result of Eq. (2), i.e., $\sigma^{*}\propto a^{-2}\propto(N_{v}/L)^{2}=(\Delta v/\kappa)^{2}$, where we recall that $\Delta v$ is the velocity difference across the vortex row. Such a quadratic dependence on $\Delta v$ is also found in the analysis of the classical Kelvin-Helmholtz instability, and keeps holding for small number $N_{v}$ of massive vortices. Remarkably, however, for large $N_{v}$ (i.e., small $a$) massive cores feature a maximum instability growth rate which is linear in $N_{v}$:
@@ 2.3 Dissipation-induced suppression of the instability
where $\bm{v}_{s,j}$ is the superfluid velocity in the neighborhood of the vortex [77, 79], and the longitudinal and transverse friction coefficients $\alpha$ and $\alpha^{\prime}$ are both typically positive. In this framework, as was shown in Ref. [46], the maximum instability growth rate (18) reads
@@ 3.1 Point-vortex model
$\mathcal{L}_{a}=\sum_{j=1}^{N_{v}}\left\{\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\dot{\theta}_{j}-\Phi_{j}-\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right\},$ (21)
@@
    $\mathcal{H}_{a}=\sum_{j=1}^{N_{v}}\left(\Phi_{j}+\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right)$ (24)
@@ 3.3 Instability of a massless vortex necklace
$\mathcal{L}_{a}^{\prime}=\sum_{j=1}^{N_{v}}\left\{\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\left(\dot{\theta}_{j}^{\prime}+\Omega\right)-\Phi_{j}-\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right\}$ (31)
@@
$\mathcal{H}_{a}\qquad\to\qquad\mathcal{H}_{a}^{\prime}=\sum_{j=1}^{N_{v}}\left(\Phi_{j}+\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right)-\Omega\sum_{j=1}^{N_{v}}\pi\hbar n_{a}(R_{2}^{2}-r_{j}^{2}).$ (32)
@@
In all cases, the most unstable mode is the one associated to $q^{*}=q_{j=N_{v}/2}=\pi/d_{v}$, where $d_{v}\approx 2\pi r_{0}/N_{v}$ is the intervortex distance, while the mode $q=0$ is such that $\sigma(q=0)=0$. The latter property is associated to the conservation of the total angular momentum, which follows from system rotational invariance. For $N_{v}\to+\infty$, the points collapse on Eq. (1), which was derived for straight vortex rows in Ref. [43]. The structure of the normal modes for $N_{v}=6$ is illustrated in Fig. 8.
@@ 4.1 Hamiltonian description
($\mathcal{L}\left(\left\{r_{j},\theta_{j}\right\},\left\{\dot{r}_{j},\dot{\theta}_{j}\right\}\right)=\sum_{j=1}^{N_{v}}\left\{\frac{M_{c}}{2}\left(\dot{r}_{j}^{2}+r_{j}^{2}\dot{\theta}_{j}^{2}\right)+\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\dot{\theta}_{j}-\Phi_{j}-\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right\}.$ (40)
@@
$\mathcal{H}=\sum_{j=1}^{N_{v}}\left\{\frac{p_{r_{j}}^{2}}{2M_{c}}+\frac{\left[p_{\theta_{j}}-\pi\hbar n_{a}\left(R_{2}^{2}-r_{j}^{2}\right)\right]^{2}}{2M_{c}r_{j}^{2}}+\Phi_{j}+\sum_{k=1}^{N_{v}}{\vphantom{\sum}}^{\prime}V_{jk}\right\}$ (41)
@@ 4.2 Core-mass-induced suppression of the instability
This relation, illustrated in Fig. 11 as a function of the number of vortices (solid lines), well captures the results of the full numerical linear-stability analysis (dots) for a vortex necklace in an annular domain.
@@ 5 Comparison with experiments
Unfortunately, it is not possible to rigorously perform the linear-stability analysis for a (either massless or massive) vortex necklace subject to frictional forces. The reason is that, in the presence of dissipation, regular vortex necklaces (see Sec. 3.2 and Sec. 4.1) no longer constitute stationary solutions of the associated dynamical systems. This circumstance prevents the application of the standard linear-stability-analysis machinery described in Sec. 3.3 and Sec. 4.2 and would require a systematic analysis of the early-time dynamics generated by the full set of equations of motion with suitable stochastic initial conditions. However, assuming that the breakdown of the necklace is only due to the Kelvin-Helmholtz mechanism, one can estimate the corresponding instability growth rate in the presence of dissipation adapting Eqs. (16)-(18), obtained for a vortex row, to the case of a vortex necklace. This can be done upon the substitution $a=2\pi r_{0}/N_{v}$ and, on the basis of what was discussed in Sec. 4.2 (and illustrated in Fig. 11), the resulting equations are expected to well capture the properties of the actual vortex necklace.
@@ 6 Conclusions
In conclusion, we have elucidated the role played by confinement, vortex mass and mutual friction on the dynamical instability of many-vortex systems. Further studies are needed to understand the role of mutual friction in the breakdown of rotating vortex necklaces. This should require the explicit solution of the full set of equations of motion with suitable stochastic initial conditions, thus providing a more reliable estimate of $\sigma^{*}$ for the case of a rotating necklace in a dissipative superfluid. Moreover, it could be interesting to analyze other effects which may affect the Kelvin-Helmholtz instability, such as non-zero temperatures, unequal vortex-core mass [83], or the coupling of vortices to sound.
@@ Acknowledgements
A. R. received funding from the European Union’s Horizon research and innovation programme under the Marie Skłodowska-Curie grant agreement Vortexons (no. 101062887). P. M. and A. R. acknowledge support by the Spanish Ministerio de Ciencia e Innovación (MCIN/AEI/10.13039/501100011033, grant PID2020-113565GB-C21), and by the Generalitat de Catalunya (grant 2021 SGR 01411). P. M. further acknowledges support by the ICREA Academia program. M. Cal., A. R. and P. M. would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) and the LabEx CARMIN (ANR-10-LABX-59-01) for their support.